Thirty ounces of pure silver is added to 50 ounces of a silver alloy that is 20% silver. What is the percent concentration of silver in the resulting alloy?
Answer by stanbon(75887)

Thirty ounces of pure silver is added to 50 ounces of a silver alloy that is 20% silver. What is the percent concentration of silver in the resulting alloy?
---------
Equation:
silver + silver = silver
30 + 0.20*50 = x(30+50)
30 + 10 = 80x
50 = 80x
x = 5/8 = 0.625 = 62.5%
===========================
Cheers,
Stan H.